# On the quantum chromatic number of Hamming and generalized Hadamard graphs

### v1

@@ Abstract
Quantum coloring finds applications in quantum cryptography and information. In this paper, we study the quantum chromatic numbers of Hamming graphs and a generalization of Hadamard graphs. We investigate the separation between the quantum and classical chromatic numbers of these graphs and determine the quantum chromatic numbers for some of them.
For the upper bounds of the quantum chromatic numbers, we develop a linear programming approach over the Hamming scheme to construct modulus-one orthogonal representations. For the lower bounds, we determine the minimum eigenvalues for some of these graphs to derive corresponding spectral lower bounds on their quantum chromatic numbers.
@@ 1 Introduction
Graph colouring plays a central role not only in combinatorics but also in quantum information theory and communication [2]. The quantum chromatic number of a graph $G$, denoted by $\chi_{Q}(G)$, was first suggested by Patrick Hayden (private communication, as reported in [2]) and independently introduced in [4]. Hadamard graphs $\varOmega_{n}$, which are defined when $n$ is a multiple of $4$, are graphs on vertex set consisting of all $\pm 1$}-vectors of length $n$, where two vertices are adjacent if and only if they are orthogonal. They can also be regarded as binary Hamming graphs with distance $n/2$, namely $H(n,2,n/2)$. These graphs provide a notable example of quantum advantage [2]: their quantum chromatic number satisfies $\chi_{Q}(\varOmega_{n})\leq n$, while the combinatorial result of Frankl and Rödl [9] implies that, for sufficiently large $n$, the classical chromatic number satisfies $\chi(\varOmega_{n})\geq(1+\varepsilon)^{n}$ for some $\varepsilon>0$, which yields an exponential separation between the quantum and classical chromatic numbers.
Despite its significance, few nontrivial lower bounds are known for the quantum chromatic number, and [10] showed that computing it is NP-hard in general. For a long time, apart from trivial classical graphs such as complete graphs, bipartite graphs, and cycles, the Hadamard graphs $\varOmega_{n}$ constituted the only known infinite family of graphs for which the quantum chromatic number could be determined.
@@
Very recently, Cao, Feng, and Tan [5] determined the quantum chromatic number of another family of binary Hamming graphs, namely H(4t1,2,2t)H(4t-1,2,2t), which forms another known infinite family of graphs with an explicitly determined quantum chromatic number. In their work, they established an upper bound on $\chi_{Q}(H(n,2,d))$ for all $d\geq\frac{n}{2}$, while leaving the case $d<\frac{n}{2}$ as an open problem.
Motivated by [5], we study the quantum chromatic number and the corresponding separation properties of general $q$-ary Hamming graphs, as well as a natural generalization of Hadamard graphs.
The $q$-ary Hamming graph of length $n$ with distance $d$, denoted by $H(n,q,d)$, is the graph on the set of $n$-tuples over a $q$-ary alphabet, where two vertices are adjacent if and only if their coordinates differ in exactly $d$ positions. The collection of all such graphs forms a symmetric association scheme known as the Hamming scheme [6].
For Hamming graphs, we develop a linear programming method over the Hamming scheme to construct orthogonal representations, whose dimensions provide upper bounds on the quantum chromatic number. Using this method, we extend the upper bound on $\chi_{Q}(H(n,2,d))$ for $d\geq\tfrac{n}{2}$ obtained in [5] to the general case $H(n,q,d)$ with $d\geq\tfrac{(q-1)n}{q}$. Moreover, the versatility of our approach allows it to handle the case $d<\tfrac{(q-1)n}{q}$ as well, where we derive additional upper bounds that address the open question posed in [5]. Our main upper bounds are summarized as follows.
Theorem 1.1.
@@
    If $\frac{(q-1)n}{q}-\frac{\sqrt{(q-1)n}}{q}<d<\frac{(q-1)n}{q}$, then $\chi_{Q}(H(n,q,d))\leq 2(q-1)^{2}\binom{n}{2}$.
@@
    If $d=\delta n$ for some $0<\delta<\frac{q-1}{q}$, then
@@
    where $h_{q}(x)=x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{1}(1-x)$ is the $q$-ary entropy function.
@@
Frankl and Rödl [9, Theorem 1.10] showed that for sufficiently large $n$ and all $d$ satisfying $\delta n<d<(1-\delta)n$ for some fixed $0<\delta<\tfrac{1}{2}$ and with $d$ even when $q=2$, there exists a constant $\varepsilon=\varepsilon(q,\delta)>0$ such that $\alpha(H(n,q,d))\leq(q-\varepsilon)^{n}$. Consequently, we have $\chi(H(n,q,d))\geq(\tfrac{q}{q-\varepsilon})^{n}$. Therefore, the first two cases in Theorem 1.1 yield an exponential separation between the quantum and classical chromatic numbers. The third case yields an MRRW-type upper bound; however, it does not lead to such a separation, which we leave as an open problem in Section 5.
In addition, we establish a Plotkin-type lower bound for the quantum chromatic number of Hamming graphs.
Theorem 1.3.
Let $n,q,d$ be positive integers with $q\geq 3$ and $d\leq n$.
@@
    If $d\geq\frac{(q-1)n+1}{q}$, then $\chi_{Q}(H(n,q,d))\geq\dfrac{qd}{qd-(q-1)n}$.
Remark 1.4.
Combining the first case in Theorem 1.1, we obtain $(q-1)(n-1)+1\leq\chi_{Q}(H(n,q,\frac{(q-1)n}{q}))\leq(q-1)n$ and $\chi_{Q}(n,q,\frac{(q-1)n+1}{q})=(q-1)n+1$. There remains a gap of (q2)(q-2) between the upper and lower bounds of $\chi_{Q}(H(n,q,\tfrac{(q-1)n}{q}))$, and we leave bridging this gap as an open question in Section 5.
We consider the following generalization of the Hadamard graph with respect to an additive group $\mathbb{G}$, denoted by $\varOmega_{n}^{(\mathbb{G})}$, which is the graph on $\mathbb{G}^{n}$ where two vertices are adjacent if and only if each element of $\mathbb{G}$ appears exactly $n/|\mathbb{G}|$ times in their difference. Note that when $\mathbb{G}$ has order 2, so it must be the cyclic group $\mathbb{Z}_{2}$, $\varOmega_{n}^{(\mathbb{Z}_{2})}$ coincides with the ordinary Hadamard graph $\varOmega_{n}$. Moreover, this generalization naturally leads to a special association scheme. The composition of an $n$-tuple $x\in\mathbb{G}^{n}$, denoted by $\operatorname{comp}(x)$, is a $|\mathbb{G}|$-tuple of nonnegative integers $(d_{g}:g\in\mathbb{G})$ such that $\sum_{g\in\mathbb{G}}d_{g}=n$. (In this paper, we use bold font to denote a composition, e.g., $\boldsymbol{d}=(d_{g}:g\in\mathbb{G})$) The composition graph of length $n$ over $\mathbb{G}$ with composition $\boldsymbol{d}=(d_{g}:g\in\mathbb{G})$, denoted by $H_{C}(n,\mathbb{G},\boldsymbol{d})$, is the directed graph on $\mathbb{G}^{n}$ in which two vertices $x$ and $y$ form a directed edge from $x$ to $y$ if and only if $\operatorname{comp}(y-x)=\boldsymbol{d}$. Indeed, all such composition graphs form an asymmetric association scheme , which we refer to as the composition scheme.  [^1] To the best of our knowledge, this scheme was first introduced by Delsarte [6] in the case $\mathbb{G}=\mathbb{F}_{q}$, who referred to it as the spectral scheme, and was later further studied by Sookoo [13] in the same setting. Note that the composition scheme with respect to an order-$q$ additive group is a refinement of the $q$-ary Hamming scheme; in particular, when $q=2$, the composition scheme coincides with the binary Hamming scheme.
For generalized Hadamard graphs, we partially determine their quantum chromatic numbers in the following two cases:
Theorem 1.5.
@@
    If $\frac{(q-1)n}{q}$ is even, then there exists $N=N(q)$ such that for every $n\geq N$,
@@
The remaining part of this paper is organized as follows. In Section 2, we introduce some basics of graph theory and present fundamental facts on the quantum chromatic number. We further discuss Hamming and composition graphs and their eigenvalues. In Section 3, we prove Theorems 1.1 and 1.3 for the quantum chromatic number of Hamming graphs. In Section 4, we prove Theorem 1.5 for generalized Hadamard graphs. Finally, in Section 5, we conclude the paper and present several open problems for future research.

### v2

@@ Abstract
As a fundamental metric for quantifying quantum advantage in non-local games, the quantum chromatic number reveals the power of entanglement in distributed tasks. In this paper, we investigate this parameter for $q$-ary Hamming graphs and a generalization of Hadamard graphs. Our main results establish an exponential separation between the quantum and classical chromatic numbers for both graph families, and determine the exact quantum chromatic numbers in several regimes. Our analysis builds on known upper and lower bounds via modulus-one orthogonal representations and minimum eigenvalues, respectively.
Previous results for Hamming graphs $H(n,q,d)$ were restricted to specific cases: the minimum eigenvalue was only identified for d>(q1)n/qd>(q-1)n/q, while modulus-one orthogonal representations had only been constructed for the binary case ($q=2$) with $d\geq n/2$. In this work, we fill several gaps in the existing literature by developing a linear programming approach to construct modulus-one orthogonal representations for arbitrary relative distances, and using the trace method to determine the minimum eigenvalues in the regime where $d$ lies slightly below the threshold (q1)n/q(q-1)n/q.
For generalized Hadamard graphs over cyclic groups and finite fields, by determining their minimum eigenvalues, we show that the spectral lower bound matches the natural upper bound on the quantum chromatic number. On the classical side, we apply the method of forbidden intersection pattern of Frankl and Rödl to obtain an exponential lower bound on the chromatic number, thereby quantifying the separation between the quantum and classical quantities.
keywords:
quantum chromatic number , Hamming scheme , generalized Hadamard graphs , linear programming , minimum eigenvalue , forbidden intersection method
journal: Journal of Combinatorial Theory, Series A
[…]
@@ 1 Introduction
Quantum entanglement enables a fascinating phenomenon known as pseudo-telepathy, where two spatially separated parties exhibit correlations so strong that they appear to possess telepathic powers. As described in the survey by Brassard, Broadbent, and Tapp [4], this phenomenon arises in non-local games where parties sharing an entangled quantum state can perform tasks that are classically impossible without communication.
In a general setting, two parties, Alice and Bob, receive inputs $(q_{A},q_{B})$ and must return answers $(a_{A},a_{B})$. A non-local game is called a pseudo-telepathy game if there exists a quantum protocol that allows them to win every instance of the game with probability $1$, whereas no classical strategy (without communication) can achieve a winning probability of $1$. This strict separation highlights the power of shared entanglement.
A canonical example of a pseudo-telepathy game is the graph coloring game [4, 10]. Formally, this game can be cast within the general framework of quantum graph homomorphisms [27], which has also been specialized to characterize quantum graph isomorphisms [2]. In the standard coloring game, a graph $G=(V,E)$ and an integer $c$ are known to both parties. A verifier sends a vertex $u$ to Alice and a vertex $v$ to Bob. Without communicating, they must return colors $c_{A}$ and $c_{B}$ from a set of $c$ colors. To win, they must satisfy the standard graph coloring constraints: they must output the same color ($c_{A}=c_{B}$) if $u=v$, and distinct colors ($c_{A}\neq c_{B}$) if $u$ and $v$ are adjacent.
Classically, Alice and Bob can win this game with probability $1$ if and only if the number of available colors $c$ is at least the classical chromatic number $\chi(G)$. If $c<\chi(G)$, any classical strategy will inevitably fail on some input pairs [5].
However, utilizing the non-local properties of quantum mechanics, players can sometimes win with strictly fewer colors. The minimum number of colors required for a perfect quantum strategy is defined as the quantum chromatic number, denoted by $\chi_{Q}(G)$ [3]. Independently, Cameron et al. [8] studied this parameter (denoted therein as $\chi_{q}(G)$) and established its rigorous mathematical foundations. Although the formal definition involves quantum states and measurement operators in Hilbert spaces (as detailed in Section 2, Definition 2.2), for the purpose of outlining our results, it suffices to understand $\chi_{Q}(G)$ as a parameter that relaxes the conditions of classical coloring.
Much effort has been devoted to the separation between classical and quantum chromatic numbers. For small graphs, Mančinska and Roberson [26] identified a graph on $14$ vertices exhibiting $\chi_{Q}(G)<\chi(G)$, which Lalonde [22] later proved to be the smallest such instance.
For infinite families, the Hadamard graphs $\varOmega_{n}$ (where $n$ is a multiple of $4$) provide a striking example of exponential quantum advantage [3]. In these graphs, the vertices correspond to $\pm 1$-vectors of length $n$, and two vertices are adjacent if and only if their vectors are orthogonal. They can be regarded as binary Hamming graphs $H(n,2,n/2)$. In 1999, Brassard et al. [5] showed that $\chi_{Q}(\varOmega_{n})\leq n$ when $n$ is a power of $2$; this result was subsequently extended to arbitrary $n$ divisible by $4$ by Avis et al. [3] in 2006. In contrast, Frankl and Rödl [17] demonstrated that the classical chromatic number is exponentially larger. Furthermore, when $n=4p^{k}$ with $p$ being a prime, the independence number $\alpha(\varOmega_{n})$, as determined by Frankl [18] and Ihringer, Tanaka [19], yield an explicit lower bound for the classical chromatic number $\chi(\varOmega_{n})$.
It is challenging to determine the exact value of $\chi_{Q}(G)$. Ji [20] proved that computing the quantum chromatic number for general graphs is NP-hard. Mančinska et al. [25] utilized the Lovász theta function to prove $\chi_{Q}(H(n,2,n/2))=n$, a result later re-derived via spectral arguments [31, 28]. For a long time, Hadamard graphs were the only non-trivial family with a known quantum chromatic number.
@@
Very recently, Cao, Feng, and Tan [9] determined the quantum chromatic number of another family of binary Hamming graphs, specifically H(4t1,2,2t)H(4t-1,2,2t), thereby identifying another infinite family with determined quantum chromatic numbers and exponential separation between quantum and classical chromatic number. In their work, they established an upper bound on $\chi_{Q}(H(n,2,d))$ for all $d\geq n/2$, while leaving the case $d<n/2$ as an open problem.
There are rare examples of graph families exhibiting determined quantum chromatic numbers and an exponential separation from the classical chromatic numbers. This scarcity motivates the further study of quantum chromatic numbers. The following research problems are of particular interest:
  • 1.
    Construct graphs where the quantum chromatic number is strictly less than the classical chromatic number. For instance, it is known that $\chi_{Q}(\varOmega_{n})=n$ when $4|n$, whereas $\chi(\varOmega_{n})\geq(1+\varepsilon)^{n}$.
  • 2.
    Determine the exact quantum chromatic numbers for specific classes of graphs.
  • 3.
    Establish tighter lower or upper bounds for the quantum chromatic numbers of various graph classes.
To tackle these problems, we rely on two fundamental bounds on the quantum chromatic number. The first is an upper bound derived from modulus-one orthogonal representations. In this paper, graphs are assumed to be simple, namely, undirected and without multi-edges and loops. An orthogonal representation of a graph $G$ of rank $k$ is a map $\rho:V(G)\to\mathbb{C}^{k}$ such that $\rho(u)$ and $\rho(v)$ are orthogonal with respect to the standard Hermitian inner product for all adjacent vertices $u,v\in V(G)$. Furthermore, a representation $\rho$ is termed modulus-one if every entry of the vector $\rho(v)$ has unit modulus for all $v\in V(G)$. Let $\xi(G)$ and $\xi^{\prime}(G)$ denote the minimum ranks of an orthogonal representation and a modulus-one orthogonal representation of $G$, respectively. Cameron et al. [8] established the following upper bound on the quantum chromatic number.
Lemma 1.1 ([8, Proposition 7]).
For any graph $G$, $\chi_{Q}(G)\leq\xi^{\prime}(G)$.
On the other hand, spectral techniques offer a lower bound. Let $\lambda_{1}\geq\dots\geq\lambda_{n}$ be the eigenvalues of the adjacency matrix of $G$. Elphick and Wocjan [16] established the following Hoffman-type bound.
Lemma 1.2 ([16]).
Let $G$ be a graph with at least one edge. Then, $\chi_{Q}(G)\geq 1+\frac{\lambda_{1}}{|\lambda_{n}|}$.
We remark that for regular graphs, applying this spectral lower bound reduces to determining the minimum eigenvalue of the graph.
1.1 Our Contributions
We investigate the quantum chromatic number and the corresponding separation properties of $q$-ary Hamming graphs, as well as a natural generalization of Hadamard graphs. Our main results establish an exponential separation between the quantum and classical chromatic numbers for both graph families, and determine the exact values in several cases.
Part I: Hamming Graphs.
The $q$-ary Hamming graph $H(n,q,d)$ is defined on the vertex set of $q$-ary $n$-tuples with adjacency determined by Hamming distance $d$. We establish tight bounds on its quantum chromatic number, making significant progress on the open question regarding the regime $d<\frac{(q-1)n}{q}$ posed in [9].
Theorem 1.3 (Upper Bounds of $\chi_{Q}(H(n,q,d))$).
@@ 1.1 Our Contributions
    If $\frac{(q-1)n}{q}>d\geq\frac{(q-1)n-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}$, then
    $\chi_{Q}(H(n,q,d))\leq\frac{\left(2(q-1)n-(q-2)\right)qd-q^{2}d^{2}}{2}.$
@@
    If $d=\delta n$ for some fixed $0<\delta<\frac{q-1}{q}$, then
@@
    where $H_{q}(x)=x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function.
@@
Building on the classical results of Frankl and Rödl [17, Theorem 1.10], which show that $\chi(H(n,q,d))$ grows exponentially for fixed relative distances, our first two cases in Theorem 1.3 establish an exponential separation between classical and quantum chromatic numbers.
The key ingredient for establishing Theorem 1.3 is a novel linear programming method over the Hamming scheme [13]. This approach provides a unified framework to efficiently construct modulus-one orthogonal representations, thereby upper bounding the quantum chromatic number.
Lemma 1.5 (Linear Programming Framework).
The quantity $\xi^{\prime}(H(n,q,d))$ is upper bounded by the value of any feasible solution to the following integer linear program:
$\begin{array}[]{lcl}\text{minimize}\quad&\sum\limits_{i=0}^{n}(q-1)^{i}\binom{n}{i}c_{i}&\\[6.0pt] \text{subject to}\quad&\begin{cases}\sum\limits_{i=0}^{n}c_{i}>0,\\ \sum\limits_{i=0}^{n}c_{i}K_{i}(d)=0,\\ c_{0},c_{1},\dots,c_{n}\in\mathbb{Z}_{\geq 0},\end{cases}\end{array}$
where $K_{i}(j)$ is the Krawtchouk polynomial given in (1). In fact, any feasible solution yields a modulus-one orthogonal representation.
To complement our upper bounds, we employ the Hoffman-type bound [16], which relies on the minimum eigenvalue of the graph. For $H(n,q,d)$, the eigenvalues are given by the Krawtchouk polynomial $K_{d}(z)$ where $z\in\{0,1,\dots,n\}$. Van Dam and Sotirov [30] conjectured that for $d\geq\frac{(q-1)n+1}{q}$, with $d$ even when $q=2$, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(1)$. Alon and Sudakov [1] proved this conjecture for $q=2$ when $n$ is large and $d/n$ is fixed. Dumer and Kapralova [15] subsequently established the result for $q=2$ and all $n$. Finally, the conjecture was proved in full generality by Brouwer et al. [6]. While the minimum eigenvalue for $d>\frac{(q-1)n}{q}$ is $K_{d}(1)$, the behavior for smaller $d$ is complex. In this work, we consider the case where $d$ lies slightly below the threshold $\frac{(q-1)n}{q}$, and we identify a second threshold, above which the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(2)$.
Lemma 1.6.
For $\frac{(q-1)n}{q}\geq d\geq\frac{(q-1)n-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-1)}+1}{q}$ with $d$ even when $q=2$, and $n$ sufficiently large, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(2)$.
Since the minimum eigenvalue of $H(n,q,d)$ is known to be $K_{d}(1)$ for $d\geq\frac{(q-1)n+1}{q}$, we can combine this fact with Lemma 1.6 to obtain the following lower bounds on $\chi_{Q}(H(n,q,d))$.
Theorem 1.7 (Lower Bounds of $\chi_{Q}(H(n,q,d))$).
Let $n,q,d$ be positive integers with $q\geq 2$, $d\leq n$, and $d$ even when $q=2$.
@@
    If $\frac{(q-1)n}{q}\geq d\geq\frac{(q-1)n-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-1)}+1}{q}$, then for sufficiently large $n$,
    $\chi_{Q}(H(n,q,d))\geq\frac{\Big(2(q-1)n-(q-2)\Big)qd-q^{2}d^{2}}{\Big(2(q-1)n-(q-2)\Big)qd-q^{2}d^{2}-(q-1)^{2}(n-1)n}.$
Comparison of Bounds. Within the range of distances where the lower bounds are established (Theorem 1.7), the ratio between the upper and lower bounds is at most $O(n)$. This gap narrows significantly for specific distances:
  • 1.
    At $d=\frac{(q-1)n+1}{q}$, the upper and lower bounds coincide, and thus $\chi_{Q}(H(n,q,d))=(q-1)n+1$.
  • 2.
    At $d=\frac{(q-1)n}{q}$, the bounds differ by an additive constant of q2q-2. Specifically,
    $(q-1)n-q+2\leq\chi_{Q}(H(n,q,d))\leq(q-1)n.$
  • 3.
    For the binary case ($q=2$) at $d=\frac{n-\sqrt{n-1}+1}{2}$, the bounds differ by a multiplicative factor $\sqrt{n-1}$. Specifically, for large $n$,
    $\frac{n^{2}-n+2\sqrt{n-1}}{2\sqrt{n-1}}\leq\chi_{Q}(H(n,2,d))\leq\frac{n^{2}-n+2\sqrt{n-1}}{2}.$
Part II: Generalized Hadamard Graphs.
Second, we extend our investigation to a natural generalization of the Hadamard graph over an abelian group $\mathbb{G}$ (written additively). We define the generalized Hadamard graph $\varOmega_{n}^{(\mathbb{G})}$ as the graph with vertex set $\mathbb{G}^{n}$, where two vertices $x,y\in\mathbb{G}^{n}$ are adjacent if and only if each element of $\mathbb{G}$ appears exactly $n/|\mathbb{G}|$ times in their difference. This construction recovers the standard Hadamard graph $\varOmega_{n}$ as the special case where $\mathbb{G}=\mathbb{Z}_{2}$.
The quantum chromatic number of generalized Hadamard graphs is always bounded above by $n$, see Lemma 4.1. If $\varOmega_{n}^{(\mathbb{G})}$ contains a clique of cardinality $n$, then we obtain a so-called generalized Hadamard matrix over $\mathbb{G}$, a concept first introduced by Drake [14]. Since the seminal work of Butson [7], numerous constructions of generalized Hadamard matrices have been developed, see for example [21], [11], and extensive surveys of diverse group structures [12]. When combined with bound $\omega(G)\leq\chi_{Q}(G)$ (see [8, Proposition 4]), these results determine $\chi_{Q}(\varOmega_{n}^{(\mathbb{G})})$ for certain choices of $\mathbb{G}$ and $n$.
In this work, we focus on generalized Hadamard graphs defined over the cyclic group $\mathbb{Z}_{q}$ and the finite field $\mathbb{F}_{q}$. By employing the trace method and character sum properties, we derive the minimum eigenvalues for $\varOmega_{n}^{(\mathbb{Z}_{q})}$ and $\varOmega_{n}^{(\mathbb{F}_{q})}$ across a wide range of parameters. Applying the Hoffman-type bound to these results allows us to establish the lower bound $\chi_{Q}\geq n$ for the quantum chromatic number. Given the natural upper bound $\chi_{Q}\leq n$, we identify new infinite families of graphs with determined quantum chromatic numbers.
Theorem 1.8 (Exact values of quantum chromatic number).
@@
    There exists a positive integer $N=N(q)$ such that for all $n\geq N$ where $\frac{(q-1)n}{q}$ is even, we have
@@
In contrast to the linear growth of the quantum chromatic number, we show that the classical chromatic number grows exponentially with $n$. For the group $\mathbb{Z}_{q}$, we construct intersection matrices and adapt the powerful method established by Frankl and Rödl [17] to prove that $\chi(\varOmega^{(\mathbb{Z}_{q})}_{n})$ is exponentially large. For the finite field $\mathbb{F}_{q}$, we derive a similar exponential lower bound by reducing the problem to the Hamming graph case.
Theorem 1.9 (Exponentially large chromatic numbers).
Let $q$ be a positive integer, and let $n$ be a positive integer divisible by $q$.
  1. 1.
    There exist positive constants $N=N(q)$ and $\varepsilon=\varepsilon(q)$ such that for all $n\geq N$ where $\frac{(q-1)n}{q}$ is even, we have
    $\chi(\varOmega_{n}^{(\mathbb{Z}_{q})})\geq(1+\varepsilon)^{n}.$
  2. 2.
    If $q$ is a prime power, there exist positive constants $N=N(q)$ and $\varepsilon=\varepsilon(q)$ such that for all $n\geq N$, with the condition that $n$ is divisible by $4$ if $q=2$,
    $\chi(\varOmega_{n}^{(\mathbb{F}_{q})})\geq(1+\varepsilon)^{n}.$
Combining Theorem 1.8 and Theorem 1.9, we confirm that these generalized Hadamard graphs also exhibit exponential quantum-classical separation.
The paper is organized as follows. In Section 2, we review the necessary background on quantum information and graph theory. Section 3 is devoted to Hamming graphs, where we establish the upper and lower bounds given in Theorems 1.3 and 1.7. We then focus on generalized Hadamard graphs: Section 4 determines their exact quantum chromatic numbers (Theorem 1.8), while Section 5 establishes the exponential lower bounds on their classical chromatic numbers (Theorem 1.9). Finally, we conclude with open problems in Section 6.